As a calculated percentage, how efficient is a synchronous data link protocol whose frames include 24 overhead bytes (frame delimiters, head fields, error check field, etc.) and a 1524-byte data field?
Answer #1

Efficiency can be calculated by the no.of information bits transmitted by the total no.of bits in the transmission.

Here no.of information bits = 1524*8 (1 byte= 8 bits) = 12192 bits
Total bits transmitted = information+ overhead bits = (1524+24)*8 = 12384

Efficiency = 12192/12384= 0.9844
% Efficiency = Efficiency X 100 = 98.44%